OpenStudy (gdgutierrez):

Find the equation of the surface that is the set of all points whose absolute value of the difference of their distances from the points (2,0,0) and (-2,0,0) equals 6 I'm having trouble understanding the question maybe because of the wording. Idk, can someone clear it up for me?

OpenStudy (ikram002p):

its a definition of something :) (Hint :- from conic )

OpenStudy (gdgutierrez):

I'm sorry, I dont understand. I thought it had to do something with the distance formula. Am I wrong?

OpenStudy (gdgutierrez):

Since there is no yz plane.

OpenStudy (dumbcow):

\[| \sqrt{(x-2)^2 +y^2 +z^2} - \sqrt{(x+2)^2 +y^2 +z^2}| = 6\] \[|u| = \sqrt{u^2}\] \[(\sqrt{(x-2)^2 +y^2 +z^2} - \sqrt{(x+2)^2 +y^2 +z^2})^2 = 36\]

OpenStudy (gdgutierrez):

Ahh, Okay I was on the right track after all but the wording would throw me off. Thank you so much!

OpenStudy (ikram002p):

another approach , this is the def of hyperbola :)

OpenStudy (gdgutierrez):

Got it! Thanks! :)